artlover03:

Heather is putting money into a savings account. She starts with 450 in the savings account, and each week she adds 60. Let S represent the total amount of money in the savings account (in dollars), and let W represent the number of weeks Heather has been adding money. Write an equation relating S to W. Then use this equation to find the total amount of money in the savings account after 13 weeks.

artlover03:

I think its 450+60w=s after 13 weeks= 1230

Pir8booty:

is it just asking for the equation

Pir8booty:

or to get the variable

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Pir8booty is it just asking for the equation \(\color{#0cbb34}{\text{End of Quote}}\) No its asking for u to solve the equation

Pir8booty:

oh ok

artlover03:

the equation and how much would it be after 13 weeks

Pir8booty:

ya the equation is right

Pir8booty:

im looking for the answer

artlover03:

I did?

jhonyy9:

correct @artlover03

Pir8booty:

its 450 +60w=s

justus:

S=60w+450

justus:

replace the w with 13 to get your answer

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 I think its 450+60w=s after 13 weeks= 1230 \(\color{#0cbb34}{\text{End of Quote}}\) here solved it correct

justus:

Yep!